[Brahms-l] Shift report 20050129 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Jan 29 2005 - 08:10:49 EST
1. Good! 5 hour store from midnight and then just one hour down time before we got the beaem back.

2. run  tr2  tr3
12655  130k 252k  
12656  120k 208k
12657   58k
12658   12k
12659   11k
12660   26k
12661   27k
12662   63k
12663  137k
12664   29k

FS@8deg A1219 MRS@45deg A1050

3. From run 12657, mainframe MRS1440 went down. See elog. We will run this weekend with only FS.


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 29 08:11:18 2005

This archive was generated by hypermail 2.1.8 : Sat Jan 29 2005 - 08:11:35 EST